Two marbles are drawn from a bag containing 6 white, 4 red, and 6 green marbles. Find the probability of both marbles being green.
Found 2 solutions by LOL MASTER, Alan3354:
Answer by LOL MASTER(6)

2/16
Two sixteenths

Answer by Alan3354(69443)

Two marbles are drawn from a bag containing 6 white, 4 red, and 6 green marbles. Find the probability of both marbles being green.
-----------
16 total marbles
---
Without replacing the 1st marble:
= (6/16)*(5/15) = 1/8
--------
with replacement:
= (6/16)^2 = 9/64